Michael is 4 years older than Tom. 5 years ago, Michael was twice tom's age? How old is Tom?
Answer by checkley77(12844)

M=X+4 NOW
T=X NOW.
5 YEARS AGO:
M=X+4-5
T=X-5
X-1=2(X-5)
X-1=2X-10
X-2X=-10+1
-X=-9
X=9 YEARS OLD. FOR TOM NOW.
9+4=13 YEARS FOR MICHAEL NOW.
PROOF:
9-1=2(9-5)
8=2*4
8=8